# Neutrino Telescope Event Classification On Quantum Computers

### v1

@@ III.1 Neural projected quantum kernels
To implement the iterative training strategy, the 2-qubit QNN is initialized as follows: the parameters for the single-qubit gates of the first qubit are set to the optimal values obtained from the 1-qubit QNN training, i.e., $\vec{\theta}^{(1)}_{i}=\vec{\theta}^{*}_{i}$ for $i=1,\dots,L$, while the parameters for the two-qubit gates are initialized to zero, i.e., $\vec{\varphi}_{i}=\vec{0}$ for $i=1,\dots,L$. The parameters for the single-qubit gates of the second qubit, $\vec{\theta}_{i}^{(2)}$, can be initialized randomly. After this initialization, the optimization is performed over all parameters. Upon completion of the training, we obtain the 2-qubit quantum embedding $\mathrm{QNN}^{(2)}_{\vec{\theta}^{*},\vec{\varphi}^{*}}(\cdot)$.
@@
This defines the kernel matrix $K$ with elements $k_{ij}=k(\vec{x}_{i},\vec{x}_{j})$. For further details on the construction and training methodology of neural quantum kernels, we refer the reader to the original work in Ref. [32]. One key aspect of this quantum kernel is that, since each kernel entry is computed using single-qubit reduced density matrices, the training phase can be less demanding in terms of quantum processing unit (QPU) usage. Instead of estimating all $M(M-1)/2$ pairwise fidelities on a quantum device—as is required in fidelity quantum kernels—only the $M$ reduced density matrices associated with the training points need to be obtained. Each of these $2\times 2$ matrices can be reconstructed from three single-qubit Pauli expectation values and stored compactly on a classical computer, where all pairwise Hilbert–Schmidt inner products can then be computed without further quantum processing.
@@ IV Results
Due to the current limitations of quantum hardware, encoding the full high-dimensional input data is infeasible. To address this, we employ physics-inspired feature engineering to reduce the data to three informative features per event. We explore two feature sets: (a) the three principal moments of inertia, $\vec{x}=(I_{2},I_{1},I_{0})$, and (b) two moment differences and the center-of-mass (CoM) traveled distance, $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$. These engineered representations, visualized in Figure 2, exhibit clear class separation, particularly in the 3D projections, thereby simplifying the classification task. Much of the model performance can thus be attributed to the discriminative power of this compact feature space, rather than to algorithmic complexity alone.
We conducted numerical experiments using the NPQK approach. A 2-qubit QNN with depth $L=10$ was used to generate the embedding. Each NPQK model was trained on 200 events and tested on 120, using the Adam optimizer with a learning rate of 0.05, a batch size of 8, and 20 training epochs. After training, the QNN served as a quantum feature map for kernel computation. The kernel matrix $K$ was evaluated under two settings: ideal simulation (purple curve) and execution on IBM’s 127-qubit Strasbourg quantum processor (blue curve). To enhance scalability, we used disjoint qubit pairs to parallelize the estimation of reduced density matrices, as described in section III.1. Each matrix was reconstructed from 1024 measurements per Pauli observable.
@@
While ideal simulation yields the highest accuracies overall, real hardware results are highly competitive. Notably, for the three highest energy bins, the test accuracy of the NPQK approach on real quantum hardware approaches 80%, closely matching ideal simulation. This is a remarkable result, especially considering the quantum circuit depth of 10 layers, which would typically be susceptible to noise and decoherence. The closeness of real and simulated performance underscores the robustness of the NPQK protocol, and in particular, validates the hybrid strategy where the quantum embedding is optimized classically, and only the quantum kernel is computed on the device. This separation allows the most noise-sensitive operations (training) to occur in a noise-free environment, while still leveraging the expressivity of quantum models during inference, thus demonstrating the practicality of current quantum resources for nontrivial classification tasks.
@@
In the low-energy regime (below 10 TeV), model performance declines, with test accuracy reaching approximately 65%, despite a training accuracy of $\sim$75%. Compared to classical machine learning algorithms within this energy range, a similar accuracy is observed: a Boosted Decision Tree (BDT) achieves an accuracy of 67%, while a GNN attains 71% [24]. It is important to emphasize that the preprocessing and data cleaning procedures employed for these classical algorithms are considerably more extensive. For instance, they include requirements such as a minimum number of non-isolated hits within the signal region, an event vertex contained in a detector volume, starting/stopping positions of events near the detector, etc.; by contrast, our approach applies only a minimal cut on the number of hits (three or more). These results indicate that, at least for events containing limited information like those found at low energies ($\leq$ 1 TeV), the performance of quantum and classical approaches is comparable. The observed equivalence can largely be attributed to the dimensionality reduction technique adopted here, which allows our classification problem to be addressed using a feasible number of qubits while still maintaining much in the way of expressibility. These findings highlight the critical role of physics-informed feature construction in quantum machine learning. Given current constraints on qubit count and circuit depth, constructing low-dimensional yet informative input representations is key to achieving strong performance. The competitive results observed on real quantum hardware underscore the value of careful data preprocessing in enabling practical quantum classification.
@@ V Conclusion
A significant success of our study lies in the formulation of a moment-of-inertia-based data encoding scheme. This physics-inspired representation enables quantum models to efficiently learn and establishes a template that can be readily extended to other datasets in neutrino astronomy.
@@ Acknowledgments
We would like to thank Sean Chisholm for their help on the development of the earlier stages of the code used in this work. We would also like to thank Brad Baxley for their help making the cascade and track event display. PZ is supported by the David & Lucile Packard Foundation. CAA are supported by the Faculty of Arts and Sciences of Harvard University, the National Science Foundation (NSF), the NSF AI Institute for Artificial Intelligence and Fundamental Interactions, the Research Corporation for Science Advancement, and the David & Lucile Packard Foundation. CAA and PZ also acknowledge kind support and enriching environment of the Harvard Quantum Initiative; in particular, in the support of undergraduate researchers involved in the earlier stages of this project. PRG has received funding from the Basque Government through the “Plan complementario de comunicación cúantica” (EXP.2022/01341) (A/20220551). MS and PRG acknowledge support from HORIZON-CL4-2022-QUANTUM01-SGA project 101113946 OpenSuperQ-Plus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, and the “Generación de Conocimiento” project Grant No. PID2021-125823NA-I00 funded by MICIU/AEI/10.13039/501100011033, by “ERDF Invest in your Future” and by FEDER EU. We also acknowledge support from the Basque Government through Grants No. IT1470-22, the Elkartek project KUBIT KK-2024/00105, and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. This work has also been partially supported by the Ministry for Digital Transformation and the Civil Service of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.

### v2

@@ III.1 Neural projected quantum kernels
[…]
@@
This defines the kernel matrix $K$ with elements $k_{ij}=k(\vec{x}_{i},\vec{x}_{j})$. For further details on the construction and training methodology of neural quantum kernels, we refer the reader to the original work in Ref. [32]. One key aspect of this quantum kernel is that, since each kernel entry is computed using single-qubit reduced density matrices, the training phase can be less demanding in terms of quantum processing unit (QPU) usage. Instead of estimating all M(M1)/2M(M-1)/2 pairwise fidelities on a quantum device—as is required in fidelity quantum kernels—only the $M$ reduced density matrices associated with the training points need to be obtained. Each of these $2\times 2$ matrices can be reconstructed from three single-qubit Pauli expectation values and stored compactly on a classical computer, where all pairwise Hilbert–Schmidt inner products can then be computed without further quantum processing.
@@ IV Results
Due to the current limitations of quantum hardware, encoding the full high-dimensional input data is infeasible. To address this, we employ physics-inspired feature engineering to reduce the data to three informative features per event. We explore two feature sets: (a) the three principal moments of inertia, $\vec{x}=(I_{2},I_{1},I_{0})$, and (b) two moment differences and the center-of-mass (CoM) traveled distance, $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$. These engineered representations, visualized in Figure 2, exhibit clear class separation, particularly in the 3D projections, thereby simplifying the classification task. Much of the model performance can thus be attributed to the discriminative power of this compact feature space, rather than to algorithmic complexity alone. This physics-inspired feature selection enables high classification performance—even in classical machine learning models—with drastically fewer features than typically required in brute-force approaches, which often demand orders of magnitude more inputs to achieve comparable accuracy; the resulting compact representation not only benefits classical methods but also facilitates the application of quantum algorithms by aligning with current hardware constraints.
We conducted numerical experiments using the NPQK approach. A 2-qubit QNN with depth $L=10$ was used to generate the embedding. Each NPQK model was trained on 200 events and tested on 120, using the Adam optimizer with a learning rate of 0.05, a batch size of 8, and 20 training epochs. After training, the QNN served as a quantum feature map for kernel computation. The kernel matrix $K$ was evaluated under two settings: ideal simulation (purple curve) and execution on IBM’s 127-qubit Strasbourg quantum processor (blue curve). To enhance scalability, we used disjoint qubit pairs to parallelize the estimation of reduced density matrices, as described in section III.1. Each matrix was reconstructed from 1024 measurements per Pauli observable. For comparison, results obtained with classical machine learning methods on the same feature sets are presented in Appendix A.
@@
While ideal simulation yields the highest accuracies overall, real hardware results are highly competitive. Notably, for the three highest energy bins, the test accuracy of the NPQK approach on real quantum hardware approaches 80%, closely matching ideal simulation. This is a remarkable result, especially considering the quantum circuit depth of 10 layers, which would typically be susceptible to noise and decoherence. The closeness of real and simulated performance underscores the robustness of the NPQK protocol, and in particular, validates the hybrid strategy where the quantum embedding is optimized classically, and only the quantum kernel is computed on the device. This separation allows the most noise-sensitive operations (training) to occur in a noise-free environment, while still leveraging the expressivity of quantum models during inference, thus demonstrating the practicality of current quantum resources for nontrivial classification tasks. A detailed analysis of hardware performance and resource budgeting is provided in Appendix B.
@@
In the low-energy regime (below 10 TeV), model performance declines, with test accuracy reaching approximately 65%, despite a training accuracy of $\sim$75%. Compared to classical machine learning algorithms within this energy range, a similar accuracy is observed: a Boosted Decision Tree (BDT) achieves an accuracy of 67%, while a GNN attains 71% [24]. It is important to emphasize that the preprocessing and data cleaning procedures employed for these classical algorithms are considerably more extensive, for instance, they include requirements such as a minimum number of non-isolated hits within the signal region, an event vertex contained in a detector volume, starting/stopping positions of events near the detector, etc.; by contrast, our approach applies only a minimal cut on the number of hits (three or more). Directly comparing quantum and classical algorithms on a fair basis, such as by fixing parameters or computational resources, is inherently challenging due to their fundamentally different natures; this difficulty mirrors comparisons even among classical models, where equitably assessing the resource demands of a convolutional neural network against those of a random forest, for instance, is nontrivial. These results indicate that, at least for events containing limited information like those found at low energies ($\leq$ 1 TeV), the performance of quantum and classical approaches is comparable. The observed equivalence can largely be attributed to the dimensionality reduction technique adopted here, which allows our classification problem to be addressed using a feasible number of qubits while still maintaining much in the way of expressibility. These findings highlight the critical role of physics-informed feature construction in quantum machine learning. Given current constraints on qubit count and circuit depth, constructing low-dimensional yet informative input representations is key to achieving strong performance. The competitive results observed on real quantum hardware underscore the value of careful data preprocessing in enabling practical quantum classification.
Additionally, we obtained results for the imbalanced class scenario, summarized in Figure 5, which shows the F1 scores for track vs. cascade classification. Since generating samples for both event types is equally costly, we trained the models on a balanced dataset and then evaluated them on a test set with a 90:10 imbalance in favor of tracks. This choice reflects realistic conditions at the lower data-processing levels in IceCube, where the atmospheric muon rate vastly exceeds the cascade rate. The experimental setup is otherwise consistent with that used for the accuracy results across different energy ranges, including the number of data points and the parallelization strategy on real hardware.
The results correspond to the energy range of 1-10 TeV and use the feature set composed of the two moment-of-inertia differences and the center of mass. We report performance for two classical algorithms—Random Forest and Support Vector Machine with Gaussian kernel (both hyperparameter-optimized, as detailed in Appendix A)—and compare them with the neural projected quantum kernel, evaluated both in ideal simulation and on real quantum hardware (IBM Strasbourg, based on the previous Eagle r3 chip, and IBM Basque Country, based on the newer Heron r2 chip).
Overall, the quantum approach remains highly competitive with the hyper-tuned classical algorithms, particularly in the test results, and the close agreement between hardware and simulation further underscores the robustness of the method.
Figure 5: Classification results (F1 score) under class imbalance. Both training and test F1 scores are shown. The training was performed on a balanced dataset, while the test set featured a 90–10 imbalance in favor of tracks. As before, 200 training points and 120 test points were used. Results correspond to the feature set composed of two moment differences and the center of mass, $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$, within the energy range of 1–10 TeV. Shown are the results for classical algorithms–Random Forest (RF) and Support Vector Classifier with Gaussian kernel (SVC)–alongside those obtained with the neural projected quantum kernel, both in ideal simulation and on real quantum hardware (IBM Strasbourg and IBM Basque Country).
@@ V Conclusion
A key achievement of our study lies in the formulation of a moment-of-inertia-based data encoding scheme. This physics-inspired representation enables quantum models to efficiently learn and establishes a template that can be readily extended to other datasets in neutrino astronomy.
@@ Acknowledgments
We would like to thank Sean Chisholm for their help on the development of the earlier stages of the code used in this work. We would also like to thank Brad Baxley for his help making the cascade and track event display. PZ is supported by the David & Lucile Packard Foundation. CAA are supported by the Faculty of Arts and Sciences of Harvard University, the National Science Foundation (NSF), the NSF AI Institute for Artificial Intelligence and Fundamental Interactions, the Research Corporation for Science Advancement, and the David & Lucile Packard Foundation. CAA and PZ also acknowledge kind support and enriching environment of the Harvard Quantum Initiative; in particular, in the support of undergraduate researchers involved in the earlier stages of this project. PRG has received funding from the Basque Government through the “Plan complementario de comunicación cúantica” (EXP.2022/01341) (A/20220551). MS and PRG acknowledge support from HORIZON-CL4-2022-QUANTUM01-SGA project 101113946 OpenSuperQ-Plus100 of the EU Flagship on Quantum Technologies, the Spanish Ramón y Cajal Grant RYC-2020-030503-I, and the “Generación de Conocimiento” project Grant No. PID2021-125823NA-I00 funded by MICIU/AEI/10.13039/501100011033, by “ERDF Invest in your Future” and by FEDER EU. We also acknowledge support from the Basque Government through Grants No. IT1470-22, the Elkartek project KUBIT KK-2024/00105, from Grant No. IT1470-22, and the Elkartek project KUBIBIT - kuantikaren berrikuntzarako ibilbide teknologikoak (ELKARTEK25/79) and from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. This work has also been partially supported by the Ministry for Digital Transformation and the Civil Service of the Spanish Government through the QUANTUM ENIA project call – Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Classical results
To benchmark the quantum models against classical approaches, we employed two widely used machine learning methods: a support vector machine (SVM) with a radial basis function (RBF) kernel and a random forest classifier. For both models, a grid search was performed over a predefined set of hyperparameters, and the optimal combination was selected using 5-fold cross-validation. The same hyperparameter search procedure was also applied to the classical results presented in the main text for the class-imbalance scenario (Figure 5), ensuring a consistent comparison between quantum and classical models.
For the SVM, the hyperparameters considered were the regularization parameter
$C\in\{0.1,1,10,100\},$ (13)
and the kernel coefficient
$\gamma\in\{\texttt{scale},\texttt{auto},0.1,1\},$ (14)
where scale sets $\gamma=\frac{1}{n_{\text{features}}\cdot\text{Var}(X)}$ and and auto sets $\gamma=\frac{1}{n_{\text{features}}},$ with $n_{\text{features}}$ being the number of features in the dataset.
For the random forest, the hyperparameters included the maximum depth of individual trees
$\text{max\_depth}\in\{2,3,4,5\},$ (15)
and the number of trees in the ensemble
$\text{n\_estimators}\in\{25,50,100,200,500\}.$ (16)
In both cases, the combination of hyperparameters yielding the highest mean cross-validation accuracy was selected, and the resulting models were subsequently evaluated on both the training and test datasets. The corresponding results are presented in Tables 1-4.
Features strategy $\vec{x}=(I_{2},I_{1},I_{0})$ $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$
Classical algorithm SVC RF SVC RF
Training accuracy $0.707\pm 0.035$ $0.794\pm 0.046$ $0.734\pm 0.040$ $0.768\pm 0.036$
Test accuracy $0.672\pm 0.052$ $0.641\pm 0.035$ $0.690\pm 0.048$ $0.678\pm 0.039$
Table 1: Training and test accuracies for different feature strategies using two classical algorithms — Support Vector Machine with Gaussian kernel (SVC) and Random Forest (RF), both with tuned hyperparameters. Results correspond to the energy regime of 100 GeV–1 TeV.
Features strategy $\vec{x}=(I_{2},I_{1},I_{0})$ $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$
Classical algorithm SVC RF SVC RF
Training accuracy $0.819\pm 0.028$ $0.886\pm 0.012$ $0.859\pm 0.023$ $0.876\pm 0.029$
Test accuracy $0.798\pm 0.036$ $0.790\pm 0.050$ $0.825\pm 0.033$ $0.825\pm 0.025$
Table 2: Training and test accuracies using classical methods corresponding to the 10 TeV–100 TeV energy regime.
Features strategy $\vec{x}=(I_{2},I_{1},I_{0})$ $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$
Classical algorithm SVC RF SVC RF
Training accuracy $0.843\pm 0.047$ $0.868\pm 0.039$ $0.871\pm 0.009$ $0.895\pm 0.028$
Test accuracy $0.833\pm 0.026$ $0.817\pm 0.016$ $0.843\pm 0.052$ $0.853\pm 0.046$
Table 3: Training and test accuracies using classical methods corresponding to the 10 TeV–100 TeV energy regime.
Features strategy $\vec{x}=(I_{2},I_{1},I_{0})$ $\vec{x}=(I_{2}-I_{1},I_{1}-I_{0},\texttt{CoM})$
Classical algorithm SVC RF SVC RF
Training accuracy $0.813\pm 0.021$ $0.865\pm 0.049$ $0.866\pm 0.009$ $0.870\pm 0.021$
Test accuracy $0.790\pm 0.038$ $0.798\pm 0.031$ $0.868\pm 0.030$ $0.873\pm 0.028$
Table 4: Training and test accuracies using classical methods corresponding to the 100 TeV–1 PeV energy regime.
Appendix B Hardware performance and resource budget
To complement the classification results, we analyzed the performance of the quantum hardware, including the impact of shot noise and readout systematics, the mean purity of the reconstructed states, and a quantitative estimate of the computational resources required for kernel construction and inference. Every single-qubit Pauli expectation value used for reconstructing the reduced density matrices was estimated using $N_{\mathrm{shots}}=1024$ measurement shots per observable. The statistical uncertainty due to shot noise for a Pauli component $m_{i}=\langle\sigma_{i}\rangle$ is bounded by $\sigma_{m_{i}}\approx\sqrt{(1-m_{i}^{2})/N_{\mathrm{shots}}}\leq 1/\sqrt{N_{\mathrm{shots}}}$, which yields $\sigma_{m_{i}}\leq 0.031$. Since the reduced density matrix is reconstructed as in Eq. (10), and the kernel is defined via the Hilbert–Schmidt product we have
$k(\vec{x}_{i},\vec{x}_{j})=\mathrm{tr}\!\bigl(\rho_{(1)}(\vec{x}_{i})\rho_{(1)}(\vec{x}_{j})\bigr)=\tfrac{1}{2}(1+\vec{m}_{i}\cdot\vec{m}_{j}),$ (17)
where $\vec{m}_{i}$ denotes the Bloch vector associated with the reduced single-qubit state $\rho_{(1)}(\vec{x}_{i})$. Then, the propagated uncertainty on a kernel element due to shot noise can be conservatively estimated as $\Delta k_{ij}\lesssim 3/(2\sqrt{N_{\mathrm{shots}}})\approx 4.7\times 10^{-2}$. In practice, this bound represents a worst-case estimate, as correlations between components and reduced Bloch vector magnitudes lead to smaller effective uncertainties.
To assess the coherence of the quantum states used in kernel construction, we measured the mean purity
$\big\langle\mathrm{tr}(\rho^{2})\big\rangle\;=\;\big\langle k(\vec{x}_{i},\vec{x}_{i})\big\rangle,$ (18)
where the expectation value $\langle\cdot\rangle$ is defined as
$\big\langle\mathrm{tr}(\rho^{2})\big\rangle\;\equiv\;\frac{1}{|X_{\mathrm{test}}|}\sum_{\vec{x}_{i}\in X_{\mathrm{test}}}\mathrm{tr}\!\left(\rho(\vec{x}_{i})^{2}\right),$ (19)
with $X_{\mathrm{test}}$ denoting the test set corresponding to the 1–10 TeV energy range. The reported value is further averaged over five random dataset splits. For single-qubit reduced states, the minimal (fully mixed) purity is $1/2$. The measured purities were $\langle\mathrm{tr}(\rho^{2})\rangle_{\mathrm{Strasbourg}}=0.80\pm 0.02$ and $\langle\mathrm{tr}(\rho^{2})\rangle_{\mathrm{Basque}}=0.85\pm 0.03$. These values indicate that, despite the 10-layer circuit depth, a significant degree of quantum coherence is maintained, and the higher purity observed on the Basque device confirms the improved performance of the Heron r2 processor compared to the Eagle r3 architecture. The observed purities are well above the fully mixed limit and ensure that the quantum kernel remains physically meaningful and not dominated by classical noise.
An important practical benefit of our PQK construction is its reduced quantum resource demand compared to fidelity-based quantum kernels. In fidelity kernels, constructing the full training kernel matrix for $N_{\mathrm{train}}$ data points typically requires $\mathcal{O}(N_{\mathrm{train}}^{2})$ circuit evaluations—specifically $N_{\mathrm{train}}(N_{\mathrm{train}}-1)/2$ pairwise fidelity estimates—while inference with $N_{\mathrm{test}}$ points involves an additional $N_{\mathrm{train}}\times N_{\mathrm{test}}$ evaluations. In our PQK scheme, by contrast, only $N_{\mathrm{train}}$ and $N_{\mathrm{test}}$ reduced density matrices need to be reconstructed on the quantum device, each from three single-qubit Pauli measurements, after which all pairwise Hilbert–Schmidt inner products are computed classically. The use of 40 disjoint qubit pairs on the IBM devices enables the simultaneous acquisition of multiple reduced density matrices, effectively reducing the number of quantum circuit executions to $N_{\mathrm{circ}}^{(\mathrm{train})}=\lceil N_{\mathrm{train}}/40\rceil$ and $N_{\mathrm{circ}}^{(\mathrm{test})}=\lceil N_{\mathrm{test}}/40\rceil$. For instance, with $N_{\mathrm{train}}=100$ and $N_{\mathrm{test}}=100$, our PQK implementation requires only three runs for training and three for inference, each measuring 40 qubit pairs in parallel. Although similar parallelization strategies could in principle be applied to fidelity-based kernels, our approach naturally incorporates this parallel structure due to the independence of the single-qubit measurements, leading to a considerably lower experimental overhead and making the method particularly well-suited to current NISQ devices.
Appendix C IceCube Systematics
Systematic uncertainties in IceCube refer to non-statistical errors that affect the measurement and interpretation of neutrino events. A significant source of systematic error comes from uncertainties in the optical properties of the ice where the detector is embedded. In particular we investigated the depth-dependent absorption and scattering lengths that govern photon propagation.
[…]